vidyo

vidyo link

Action Items:

News:

  • from last CB:
    • all nominees were approved -> congrats to Martino M. and Paul
    • timing layer approved -> role of padova about mechanics (Bellettoni)

Framework e produzione

  • no special news. few additional datasets trasferred to legnaro
  • meeting about cloud cluster for deep learning tools yesterday -> progress on-going
    • Paul is working on it - DianaHep tools on spark.
    • studente triannale is also going to use open access code from h-tautau analysis based on these tools.
  • CLEANING OF LUSTRE -> space needed. Everybody must check his area
  • problem with farm in Legnaro last week -> cooling system. Nodes in Legnaro are now fine
  • lsf queues -> there is a skype chat to manage the usage of lsf. everybody who wants to use them can be added to the group

Attivita' di analisi

B physics

  • trigger -> proposal for new menu presented
  • finishing tests based on some feedback from Moriond, before CWR:
    • check for contamination from intermediate states of di-mesons. Probably not an issue in CMS due to low rate of fake muons.

Dark Matter

  • no news. we need to see if Siew will come to Padova.

Dibosoni, VZ risonante (V=W,Z)

invisible:
  • presentation last week.
  • converging to pre-approval asap but Easter break and ifae will probably slow it down
  • analysis is ready. checks from convenors done. paper is also okay and will be circulated. lept channel is on jacopo's shoulders and decoupled from Lisa's channel.

hh4b non risonante

  • analysis is almost complete. just trigger efficiency study is missing.
  • status presented two weeks ago to hbb
  • pre-approval requested -> probably Wed 19th
  • target is LHCP but also

Same sign leptons

  • di-mu and mu-ele channels done.
  • final version of the note will come just after Easter. Presentation will be required to convenors then.
  • 3 independent methods for irred. bkg evaluation. Good results

Seesaw a 13 TeV

  • PAS made. work is on-going to get paper (Carla will be added to author list)
  • for 13TeV a new strategy was adopted to get th interpretation (fig.13 of pas).
  • generalization of scenario applied now and new figure has been released.

EPR

  • martino D. did 2 weeks of shifts for Pixel commissioning (EPR core)
  • patrizia signed for offline run manager

Isolamento muoni

Flavor tagging

Trigger

Detector

pixel phase1
  • new pixel detector has been inserted. cooling is on.
  • some delays due to cooling and calibration tasks, especially for BPIX.
  • modules are working fine but final percentage will not be 100% due to issues with readout chain (dc-dc converters, FEDs,..).

A.O.B.:

  • created spreadsheet to keep track of students' activities/thesis.
  • 2 students will go at Cern with 'trainee' contract.
  • Gaia Grosso is working on ttH with sparlk tools
  • student from Astronomy (start in summer)
  • another one at cern to work for future collider
  • Matteo Migliorini is working on ttH (target Sept)
  • Valeria Fioroni is working on HH (target Sept)
  • Giovanni is going to be summer student

-- MarcoZanetti - 2016-02-01

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2017-04-11 - MartinoDallOsso
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback